A wire of length 2.5 km and resistance 35 ohm has fallen from a height of 10m in earth's horizontal field of 2 *10^-5 T . What is the current through the wire?
 
Please provide the formulae used and solution..
please!

Grade:12

2 Answers

Vikas TU
7 years ago
I think this should go as:
 
potential generation = E*d = 2*10^-5*10 = 2*10^-4 V.
I = V/Resistance
  =  2*10^-4/35 A
is the required answer.
Perhaps!
Nisha singh
4 years ago
A wire of length 2.5km in part of circuit of resistance 35ohm oriented along E-W direction falls under